OpenStudy (anonymous):

slope=-2,through (-7,5)

OpenStudy (anonymous):

Guessing you want this in y = mx+b form?

OpenStudy (muzzack):

Okay so lets replace (x,y) in the equation y-y1=m(x-x1)

OpenStudy (muzzack):

i got this everyone

OpenStudy (anonymous):

:)

OpenStudy (muzzack):

so lets replace y and x and m now that they have given you the slope y-5=-2(x+7)

OpenStudy (muzzack):

solve it: y-5=-2x-14 y=-2x-9 there you go ;)

OpenStudy (muzzack):

now graph it